OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

OpenStudy (anonymous):

convert it to y = mx + b form, then m is slope and b is y-intercept

OpenStudy (anonymous):

an u give me an example

OpenStudy (anonymous):

suppose you have 3x - 4y = 13 then you rearrange 3x - 13 = 4y further 3x - 13 y = --------- 4 so, 3 13 y = ---- x - ----- 4 4 13 so m = 3/4 and y-intercept = - ----- 4

OpenStudy (anonymous):

awesome thank you:)

OpenStudy (anonymous):

u r welcome....☺